### v1

# LLM-Powered Social Digital Twins: A Framework for Simulating Population Behavioral Response to Policy Interventions

Abstract
Predicting how populations respond to policy interventions is a fundamental challenge in computational social science and public policy. Traditional approaches rely on aggregate statistical models that capture historical correlations but lack mechanistic interpretability and struggle with novel policy scenarios. We present a general framework for constructing Social Digital Twins virtual population replicas where Large Language Models (LLMs) serve as cognitive engines for individual agents. Each agent, characterized by demographic and psychographic attributes, receives policy signals and outputs multi-dimensional behavioral probability vectors. A calibration layer maps aggregated agent responses to observable population-level metrics, enabling validation against real-world data and deployment for counterfactual policy analysis.
We instantiate this framework in the domain of pandemic response, using COVID-19 as a case study with rich observational data. On a held-out test period, our calibrated digital twin achieves a 20.7% improvement in macro-averaged prediction error over gradient boosting baselines across six behavioral categories. Counterfactual experiments demonstrate monotonic and bounded responses to policy variations, establishing behavioral plausibility. The framework is domain-agnostic: the same architecture applies to transportation policy, economic interventions, environmental regulations, or any setting where policy affects population behavior. We discuss implications for policy simulation, limitations of the approach, and directions for extending LLM-based digital twins beyond pandemic response.
1 Introduction
Governments continuously face the challenge of predicting how populations will respond to proposed policies. Will a carbon tax reduce driving, or will people simply absorb the cost? How will citizens react to a new social welfare program? Will businesses comply with environmental regulations? These questions are fundamentally about behavioral simulation modeling how heterogeneous individuals make decisions in response to changing incentives and constraints.
Traditional approaches to policy prediction fall into two categories. Aggregate statistical models (econometric regressions, time series forecasting) capture historical correlations but lack mechanistic interpretability: they can say what happened but not why [1]. When policies are novel or contexts shift, these models extrapolate poorly. Agent-based models (ABMs) offer mechanistic richness by simulating individual decision-makers [2, 3], but require extensive manual specification of decision rules a knowledge bottleneck that limits applicability to domains where behavior is well-understood.
Recent advances in Large Language Models (LLMs) suggest a new paradigm. LLMs trained on vast corpora of human-generated text have learned implicit models of human reasoning, preferences, and decision-making [4, 5]. Researchers have demonstrated that LLMs can reproduce survey responses when conditioned on demographic attributes [6], engage in realistic social interactions [7], and serve as proxies for human subjects in economic experiments [8]. These capabilities suggest using LLMs as cognitive engines for agent-based simulation replacing hand-coded decision rules with neural models that have learned human-like reasoning.
1.1 The Digital Twin Paradigm
We propose Social Digital Twins: virtual replicas of populations where each simulated individual is powered by an LLM conditioned on demographic and contextual attributes. The framework consists of four components:
  1. 1.
    Agent Population: A set of synthetic personas representing the target population’s demographic distribution
  2. 2.
    LLM Cognitive Engine: A language model that generates behavioral outputs given persona attributes and policy context
  3. 3.
    Calibration Layer: A learned mapping from agent probability outputs to observable population metrics
  4. 4.
    Validation Protocol: Comparison against real-world observational data with strict temporal separation
This architecture is domain-agnostic. The same framework can simulate:
  • Pandemic response: Mobility changes under lockdown policies (our case study)
  • Transportation: Mode choice under congestion pricing
  • Environmental behavior: Consumption changes under carbon taxes
  • Financial decisions: Savings rates under interest rate changes
  • Health behaviors: Vaccination uptake under public health campaigns
The key insight is that LLMs have learned general models of human decision-making from training data spanning all these domains. By conditioning on appropriate context, we can instantiate domain-specific behavioral simulators without domain-specific training.
1.2 Case Study: COVID-19 Pandemic Response
We validate this framework using COVID-19 pandemic response data, chosen for several reasons:
  1. 1.
    Rich observational data: Google Mobility Reports [10] and survey trackers [11] provide high-frequency behavioral measurements
  2. 2.
    Policy variation: The Oxford COVID-19 Government Response Tracker [12] documents daily policy changes across 180+ countries
  3. 3.
    Global natural experiment: Unprecedented policy variation enables studying behavioral responses across contexts
  4. 4.
    Multi-dimensional behavior: Six mobility categories (work, retail, grocery, transit, parks, residential) capture diverse behavioral adaptations
While our evaluation focuses on pandemic response, the methodology generalizes. We emphasize that COVID-19 serves as a validation dataset, not the primary contribution. The contribution is the framework itself.
1.3 Contributions
This paper makes four contributions:
  1. 1.
    A General Framework: We formalize Social Digital Twins as LLM-powered agent simulations with calibrated output mappings, applicable across policy domains
  2. 2.
    Multi-Dimensional Behavioral Modeling: Unlike prior work extracting scalar compliance probabilities, our agents output multi-dimensional behavioral vectors enabling richer validation
  3. 3.
    Calibration Methodology: We present a multi-objective optimization approach for learning agent-to-observation mappings from historical data
  4. 4.
    Rigorous Evaluation Protocol: We demonstrate strict train/test separation, per-category metrics, counterfactual sanity checks, and ablation studies establishing component necessity
The remainder of this paper is organized as follows: Section 2 reviews related work across agent-based modeling, LLM simulation, and digital twins. Section 3 presents the general framework. Section 4 details our pandemic response instantiation and experimental results. Section 5 discusses implications, limitations, and extensions. Section 6 concludes.
2 Related Work
2.1 Agent-Based Modeling for Policy
Agent-based models (ABMs) simulate populations as collections of autonomous agents following decision rules [2]. In economics, ABMs model market dynamics [24]; in epidemiology, disease transmission [25]; in urban planning, transportation choices [26]. Classical ABMs require explicit rule specification, typically based on utility maximization [27] or behavioral heuristics [28]. This knowledge bottleneck limits applicability: modelers must anticipate relevant decision factors and encode them explicitly.
2.2 LLMs as Behavioral Simulators
Recent work explores LLMs as implicit models of human behavior. Argyle et al. [6] demonstrate “silicon sampling” LLMs conditioned on demographics reproduce political survey distributions. Park et al. [7] create “generative agents” that exhibit emergent social behaviors in simulated environments. Horton [8] shows LLMs behave reasonably in economic games, proposing them as “homo silicus” for experiments. Aher et al. [9] use LLMs to simulate human subject experiments, finding reasonable alignment on classic paradigms.
These works establish that LLMs contain useful behavioral priors. However, most lack calibration against observational data they evaluate face validity or alignment with controlled experiments rather than real-world prediction accuracy.
2.3 Digital Twins
The digital twin concept originated in manufacturing as virtual replicas of physical systems [13]. Extensions to urban systems [14], healthcare [15], and social systems [16] have been proposed. Policy-oriented digital twins aim to simulate citizen responses to interventions, but most implementations use simplified behavioral models (e.g., gravity models for mobility, discrete choice for mode selection) rather than LLM-driven reasoning.
2.4 Retrieval and Grounding for LLMs
Parallel work on improving LLM factuality through retrieval-augmented generation [17], knowledge editing [18], and attention biasing [19] is conceptually related. These methods ground LLM outputs in external knowledge; our calibration layer similarly grounds agent outputs in observational data. The key difference is that we ground behavioral predictions rather than factual claims.
2.5 Positioning This Work
Our framework combines:
  • LLM behavioral priors (from generative agent work)
  • Empirical calibration (from econometrics)
  • Multi-agent simulation (from ABM tradition)
  • Observational validation (from causal inference methodology)
Unlike prior LLM simulation work, we calibrate against real-world observational data with strict temporal holdout. Unlike traditional ABMs, we replace hand-coded rules with LLM-generated behavioral outputs. Unlike aggregate forecasting, we maintain individual-level heterogeneity and can perform counterfactual analysis.
3 Method: Social Digital Twins
3.1 Framework Overview
[…]
$\bullet$ Pandemic: Stringency
$\bullet$ Transport: Pricing
$\bullet$ Climate: Carbon tax
Any LLM:
GPT-4, Gemini,
Llama, Claude…
Figure 1: Social Digital Twin Framework. The architecture is domain-agnostic: policy signals, agent attributes, and observable outputs are instantiated per-domain, while the LLM engine and calibration methodology are general.
3.2 Agent Population
We represent the target population as a set of $N$ synthetic personas. Each persona $i$ is characterized by an attribute vector $\mathbf{a}_{i}\in\mathcal{A}$ containing:
  • Demographics: Age, gender, nationality, location
  • Socioeconomics: Occupation, income level, education
  • Psychographics: Risk tolerance, trust in institutions, social values
Attribute distributions are drawn from census data, surveys, or domain knowledge to match the target population. The persona set should capture relevant behavioral heterogeneity attributes that might influence policy response.
3.3 LLM Cognitive Engine
For each (persona, context) pair, we construct a prompt $\mathbf{x}_{i,t}$ containing:
  1. 1.
    Persona attributes $\mathbf{a}_{i}$
  2. 2.
    Current context (date, policy state, environmental factors)
  3. 3.
    Task specification requesting behavioral probability outputs
The LLM generates a response parsed into a $d$-dimensional probability vector:
$\mathbf{p}_{i,t}=\text{LLM}(\mathbf{x}_{i,t})\in[0,1]^{d}$ (1)
Each dimension represents the probability of engaging in a specific behavior (e.g., going to work, using public transit, wearing a mask). The multi-dimensional output is crucial: it enables validation against diverse observational signals and captures behavioral trade-offs (e.g., less transit use $\rightarrow$ more car use).
3.4 Aggregation
Individual agent outputs are aggregated to population-level predictions. For simple cases, we use the mean:
$\bar{\mathbf{p}}_{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{p}_{i,t}$ (2)
More sophisticated aggregations can weight by demographic prevalence or apply post-stratification to match known population totals.
3.5 Calibration Layer
Raw LLM probabilities are not directly comparable to observational metrics (which may be percentages, counts, indices, etc.). We learn a calibration function:
$\hat{\mathbf{y}}_{t}=f(\bar{\mathbf{p}}_{t};\theta)$ (3)
In our implementation, we use per-dimension linear mappings with clipping:
$\hat{y}_{t,k}=\text{clip}(\alpha_{k}\cdot\bar{p}_{t,k}+\beta_{k},y_{\min},y_{\max})$ (4)
Parameters $\theta=\{(\alpha_{k},\beta_{k})\}_{k=1}^{d}$ are learned by minimizing prediction error on training data:
$\theta^{*}=\arg\min_{\theta}\mathcal{L}(\hat{\mathbf{y}},\mathbf{y};\theta)$ (5)
We use Optuna [20] with Tree-structured Parzen Estimator for multi-objective optimization, balancing errors across behavioral dimensions.
3.6 Validation Protocol
Rigorous validation requires:
  1. 1.
    Temporal Separation: Strict train/validation/test splits preventing information leakage
  2. 2.
    Per-Dimension Metrics: Report accuracy for each behavioral category, not just aggregate
  3. 3.
    Baseline Comparison: Compare against non-agentic alternatives (statistical models, persistence)
  4. 4.
    Counterfactual Sanity: Verify monotonic, bounded responses to policy variations
  5. 5.
    Ablation Studies: Demonstrate component necessity (calibration, persona diversity, etc.)
3.7 Counterfactual Analysis
A key advantage of mechanistic simulation is counterfactual analysis. Given a trained digital twin, we can simulate alternative policy scenarios:
$\Delta\mathbf{y}=f(\text{LLM}(\mathbf{x}|\text{policy}^{\prime})))-f(\text{LLM}(\mathbf{x}|\text{policy}))$ (6)
Unlike black-box forecasting, we can ask: “What would behavior have been under different policies?” Such counterfactuals are not causally identified from observational data alone [21], but they provide behavioral plausibility checks. If the twin shows reasonable responses to policy variations (e.g., stricter policy $\rightarrow$ more compliance), we gain confidence in its internal reasoning.
4 Case Study: Pandemic Response
We instantiate the framework for COVID-19 pandemic response in the UAE.
4.1 Domain-Specific Configuration
4.1.1 Policy Signal
The Oxford COVID-19 Government Response Tracker [12] provides:
  • StringencyIndex (0–100): Composite of containment/closure policies
  • GovernmentResponseIndex: Broader measure including health/economic policies
4.1.2 Behavioral Dimensions ($d=6$)
Agent outputs map to Google Mobility categories [10]:
  1. 1.
    Retail & Recreation
  2. 2.
    Grocery & Pharmacy
  3. 3.
    Parks
  4. 4.
    Transit Stations
  5. 5.
    Workplaces
  6. 6.
    Residential
4.1.3 Observable Metrics
Google Mobility Reports provide daily percentage changes from pre-pandemic baseline (median values from January 3–February 6, 2020).
4.1.4 Agent Population
We generate $N=10$ personas reflecting UAE demographics [22, 23]:
  • Nationality: 10% UAE National, 90% Expatriate
  • Occupation: Distribution across sectors (Construction, Services, Professional, etc.)
  • Risk Perception: Low / Medium / High
4.1.5 LLM Engine
We use Gemini 2.0 Flash Lite for cost-efficiency. Prompts include persona attributes, current date, and policy stringency, requesting JSON-formatted probability outputs.
4.2 Experimental Setup
4.2.1 Data Splits
  • Train (Calibration): April 1, 2020 – March 31, 2021
  • Validation: April 1, 2021 – September 30, 2021
  • Test: October 1, 2021 – End of data availability
4.2.2 Baselines
  • Persistence: $\hat{y}_{t+1}=y_{t}$
  • Gradient Boosting (GBM): HistGradientBoostingRegressor with policy lags (0, 7, 14, 21, 28 days) and temporal features
4.3 Results
4.3.1 Main Results
Table 1: Test Set Performance (RMSE, lower is better)
Category GBM Baseline Digital Twin $\Delta$%
Retail & Recreation 23.17 14.42 +37.8%
Grocery & Pharmacy 28.31 45.67 -61.3%
Parks 28.05 18.91 +32.6%
Transit Stations 38.72 52.55 -35.7%
Workplaces 70.45 7.55 +89.3%
Residential 6.13 15.41 -151.4%
Macro Average 32.47 25.75 +20.7%
4.3.2 Analysis
The digital twin achieves 20.7% improvement in macro-averaged RMSE. Performance varies by category:
Where LLM reasoning helps:
  • Workplaces (+89.3%): LLM captures policy-work relationship (“lockdown $\rightarrow$ work from home”)
  • Retail (+37.8%), Parks (+32.6%): Discretionary activities are policy-sensitive; LLM reasoning outperforms statistical extrapolation
Where statistical baselines excel:
  • Residential (-151%): Low-variance, inertial patterns favor autoregressive features
  • Grocery (-61%), Transit (-36%): Essential/infrastructure-driven activities show less policy sensitivity
This pattern suggests LLM-based simulation adds value for decision-driven behaviors where policy semantics matter, while statistical models suffice for inertial behaviors driven by routine.
4.3.3 Counterfactual Sanity
We simulate policy variations on April 15, 2020 (peak pandemic):
Table 2: Counterfactual Policy Shocks
Scenario Stringency Stay Home % Verdict
Relaxed 60 73% Monotonic $\downarrow$
Baseline (Actual) 90 93% Historical Match
Extreme 100 95% Bounded Response
The twin demonstrates: (1) Monotonicity: Higher stringency $\rightarrow$ higher compliance, (2) Boundedness: Diminishing returns at extremes, (3) Plausibility: Responses align with intuition about pandemic behavior.
4.3.4 Ablation Studies
Table 3: Ablation Study (Macro-RMSE)
Configuration Macro-RMSE
Full System 25.75
- No calibration (raw probs) 78.32
- No clipping 26.14
- Single slope for all categories 31.42
- Uniform personas 29.87
- Single persona (N=1) 32.15
Key findings:
  • Calibration is essential (78.32 $\rightarrow$ 25.75 with calibration)
  • Per-category calibration matters (31.42 $\rightarrow$ 25.75 with separate $\alpha,\beta$)
  • Persona diversity helps (32.15 $\rightarrow$ 25.75 with 10 vs. 1 persona)
5 Discussion
5.1 Implications for Policy Simulation
Our results suggest LLM-powered digital twins are viable for policy simulation, with specific strengths and limitations:
Strengths:
  • Semantic understanding: LLMs capture policy-behavior relationships that statistical models miss
  • Generalization: Same architecture applies across policy domains with minimal modification
  • Interpretability: Agent reasoning can be inspected (unlike black-box forecasts)
  • Counterfactual analysis: Enables “what if” policy exploration
Limitations:
  • Inertial behaviors: LLMs without memory struggle with routine-driven patterns
  • Calibration dependence: Performance relies on quality of calibration data
  • Computational cost: LLM inference is more expensive than statistical models
  • Temporal grounding: LLM knowledge cutoffs may introduce anachronisms
5.2 Extensions Beyond Pandemic Response
The framework applies directly to other domains:
Transportation Policy:
  • Policy signal: Congestion pricing, parking fees, transit subsidies
  • Behavioral dimensions: Mode choice (car, transit, bike, walk)
  • Observable metrics: Traffic counts, mode share surveys
Environmental Policy:
  • Policy signal: Carbon tax, efficiency standards, recycling mandates
  • Behavioral dimensions: Energy consumption, transportation choices, waste behavior
  • Observable metrics: Utility data, emissions inventories
Economic Policy:
  • Policy signal: Interest rates, tax changes, stimulus payments
  • Behavioral dimensions: Saving, spending, labor supply
  • Observable metrics: Consumer spending data, labor force surveys
5.3 Limitations and Future Work
5.3.1 Sample Size
Our test evaluation is limited by simulation cost (10 personas $\times$ 10 dates). Production deployment requires larger populations and more comprehensive temporal coverage.
5.3.2 Inertia Modeling
Adding autoregressive components (“yesterday’s behavior influences today”) would improve predictions for routine-driven categories like residential.
5.3.3 Multi-Signal Calibration
Incorporating additional data sources (surveys, administrative records) enables richer calibration and validation.
5.3.4 Causal Identification
Our counterfactuals establish behavioral plausibility, not causal effects. Integration with causal inference methods [21] is an important direction.
5.3.5 Robustness
Repeating experiments with persona resampling and different LLMs would strengthen claims about framework generality.
6 Conclusion
We present a general framework for Social Digital Twins LLM-powered simulations of population behavioral response to policy interventions. The framework comprises synthetic agent populations with demographic attributes, LLM cognitive engines generating multi-dimensional behavioral outputs, calibration layers grounding predictions in observational data, and rigorous validation protocols.
Validating on COVID-19 pandemic response as a case study, our calibrated digital twin achieves 20.7% improvement over gradient boosting baselines on macro-averaged mobility prediction. The twin demonstrates particular strength on policy-sensitive, decision-driven behaviors (workplaces, retail) while statistical baselines excel on inertial patterns (residential). Counterfactual experiments confirm monotonic, bounded responses to policy variations.
The framework is domain-agnostic. COVID-19 provides rich validation data, but the same architecture applies to transportation, environmental, economic, or any policy domain where interventions affect population behavior. We hope this work advances the use of LLMs as cognitive engines for computational social science and policy analysis.
Data Availability
Appendix A Prompt Template
Domain-agnostic template structure:
You are simulating a specific {DOMAIN_CONTEXT}.
**Your Persona:**
- {ATTRIBUTE_1}: {value}
- {ATTRIBUTE_2}: {value}
**Current Situation:**
- {CONTEXT_1}: {value}
- {POLICY_SIGNAL}: {value}
**Task:**
Estimate your probability (0.0 to 1.0) of:
- {BEHAVIOR_1}
- {BEHAVIOR_2}
Return ONLY a JSON object.
Appendix B Pandemic Case Study: Specific Prompt
You are simulating a specific resident in the UAE
during the COVID-19 pandemic.
**Your Persona:**
- Nationality: {nationality}
- Occupation: {employment}
- Risk Perception: {risk_perception}
- Income Level: {income}
**Current Situation:**
- Date: {date}
- Policy Stringency (0-100): {stringency}
**Task:**
Estimate probability (0.0 to 1.0) of:
- "go_work_prob"
- "discretionary_outings_prob"
- "essentials_prob"
- "transit_use_prob"
- "outdoor_leisure_prob"
- "stay_home_prob"
Return ONLY a JSON object.